OpenStudy (samsan9):

how to find the tan 60 using the double angle formula?

OpenStudy (jmark):

We have double angle formula for tanx: tan(2x) = 2tanx/(1-tan2x) = 2tan60/(1-tan2(60)) = 3.464/2 = 1.732 tan 60 = √3 Icse Question Papers - http://icse.edurite.com/icse-question-papers.html